OpenStudy (anonymous):

2.5*10^5/7.6*10^6 total family net worth is......

OpenStudy (anonymous):

how many dollars

OpenStudy (anonymous):

\[\frac{ 2.5 \times 10 \times 10^4 }{ 7.6 \times 10 \times 10^5 }\] solve this now;)

OpenStudy (anonymous):

now when i did my equation i got .328947 so i rounded it to .33 then i did the .33 times 10^-2 so then it asks for the amount of dollars what do i put 33 or .0033

OpenStudy (anonymous):

so i multiply instead of divide 2.5 and 7.6

OpenStudy (anonymous):

Oops, I'm an idiot. You're totally right :-)

OpenStudy (anonymous):

but would it be .0033 dollars

OpenStudy (anonymous):

\[\frac{2.5\times10^5}{7.6\times10^6}=\frac{25}{76}\times10^{-1}\simeq3.3\times10^{-2}\]

OpenStudy (anonymous):

Does it expect a dollar? I find it tough to believe a family's net worth is a fraction of a penny...

OpenStudy (anonymous):

so would it be 0

OpenStudy (anonymous):

and it suppose to e 10^7 instead of 10^6